16. You have just received $185,000 from the estate of a long-lost rich uncle. If you invest all your inheritance in a tax-free bond fund earning 6.8% compounded quarterly, how long do you have to wait to become a millionaire? (Round your answer to two decimal places.)

20.In the following ordinary annuity, the interest is compounded with each payment, and the payment is made at the end of the compounding period.
Find the accumulated amount of the annuity. (Round your answer to the nearest cent.)

$1000 monthly at 4.2% for 20 years.

In the following ordinary annuity, the interest is compounded with each payment, and the payment is made at the end of the compounding period.

21. Find the required payment for the sinking fund. (Round your answer to the nearest cent.)

Monthly deposits earning 6% to accumulate $9000
Answer #1

16.

Number of years to become a millionare can be found by using the NPER function in excel

=NPER(0.068/4,,-185000,1000000,)/4

=25.025 years or 25.03 years

20.

We have to calculate future value of annuity, which can be calculated using the FV function,

There are 20 years, 12 months each, hence 240 months.

=FV(0.042/12,240,-1000,,)

=$375,134.95

21.

Time period is missing from the question posted, i am assuming 10 years, please comment if its something else or you can change the excel formula by substituting 120 with (12 months x no of years). Thank You.

We have to calculate monthly payment, it can be calculated using the PMT function in excel,

=PMT(0.06/12,120,,9000,)

=$54.92